OpenStudy (anonymous):

need some help with a question, it is sigma, from 1 to infinity, 3n/(3^sqrt n^4)

OpenStudy (bobo-i-bo):

Is this the question: \[\sum_{n=1}^{\infty}\frac{3n}{\sqrt{3}n^4}\] ?

OpenStudy (anonymous):

the bottom is the 3 sqrt of n^4

OpenStudy (anonymous):

was not thinking

OpenStudy (bobo-i-bo):

So this: \[\sum_{n=1}^{\infty}\frac{ 3n }{ 3\sqrt{n^4} }\] ?

OpenStudy (bobo-i-bo):

oooh

OpenStudy (bobo-i-bo):

i get it now :P

OpenStudy (bobo-i-bo):

\[\sum_{n=1}^{\infty}\frac{ 3n }{ \sqrt[3]{n^4} }\]

OpenStudy (bobo-i-bo):

Okay, so do you know how to manipulate indices?

OpenStudy (anonymous):

yeah lol, i just need somw help in seeing which method for a infinite series, like alternating series, ratio method, etc.

OpenStudy (bobo-i-bo):

Haha, sorry, kk :P

OpenStudy (bobo-i-bo):

Let me think ^_^

OpenStudy (anonymous):

thank you

OpenStudy (bobo-i-bo):

Comparison test

OpenStudy (anonymous):

awesome thanks